OpenStudy (littleheron):

will medal Calculate the slope of the line given the points (-2, 6) and (0, 4). A. -1 B. 1 C. 0 D. none of the above

OpenStudy (anonymous):

do you know how to find a slope?

OpenStudy (lolacole12):

where is the line

OpenStudy (littleheron):

no

OpenStudy (maddy1251):

Slope is rise over run.

OpenStudy (maddy1251):

\[\frac{ Rise }{ Run }\]

OpenStudy (maddy1251):

Our "Rise" is our y-coordinates. Our "Run" is our x-coordinates.

OpenStudy (maddy1251):

The formula for slope is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

she got it do you understand that ?

OpenStudy (maddy1251):

(-2, 6) and (0, 4) So your first pair of coordinates can be (-2,6), if you'd like. So. x1=-2 y1=6

OpenStudy (maddy1251):

Second pair can be (0,4), if you'd like. So, x2=0 y2=4.

OpenStudy (maddy1251):

I'll plug these in for you, but you will simplify.

OpenStudy (maddy1251):

\[\frac{ 4-6 }{ 0-(-2) }\]

OpenStudy (maddy1251):

@AlexisT20 Thank ya.

OpenStudy (littleheron):

I gtg to a classconnect, bye!

OpenStudy (maddy1251):

If there is something you don't understand, always mention me. Have a good day!

OpenStudy (anonymous):

no problem :D and bye